OpenStudy (anonymous):

If the base of a parallelogram is 30 mm and its area is 150 mm², then the height of the parallelogram is _____ mm

OpenStudy (saifoo.khan):

Hi Davette.

OpenStudy (saifoo.khan):

Area = b * h 150 = 30 * h h = 150/30 h = 5mm.